OpenStudy (kiley_justine):

What's the greatest common factor of the monomials 18x2, 27x4, 30x3

OpenStudy (trojanpoem):

18 x 2 = 2* 3 * 3 * 2 27 x 4 = 2 * 3 * 3 * 3* 2 30 x 3 = 2 * 3 * 3 * 3 so 2 * 3 * 3 = 18